HIGH RESOLUTION FTIR STUDIES OF THE ν_{2} BAND OF CH_{2}F_{2}

\underline{M.N. DEO} and K. KAWAGUCHI, Nobeyama Radio Observatory, Minamimaki, Minamisaku, Nagano, 384-1305, Japan.

The very weak ν_{2} (1509.1 cm^{−1}) band of the CH_{2}F_{2} molecule was observed for the first time in the infrared. The high resolution (0.002 cm^{−1}) FTIR absorption data of this band have been analyzed by taking into account the Coriolis interaction with the ν_{8} (1435.6 cm^{−1}) level. A strong intensity perturbation was observed, which enhances the intensity of the ν_{2} band. More than 1000 ( J”=40, Ka”=20) transitions were assigned in the ν_{2} band and a simultaneous least squares fit of the data of the ν_{2} and the ν_{8} bands was carried out. The molecular parameters and the interaction parameters obtained from the fit, satisfactorily reproduce the observed relative inetensities of the spectrum. By simulating the spectrum, we could also determine the relative sign of the transition dipole moments for ν_{2} ← 0 and ν_{8} ← 0 transitions.